A Bibliography of Publications about the Go programming language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 December 2023 Version 0.08

Title word cross-reference

action [Ken16]. Adding [MKSS18]. Advanced [Tor19]. agnostic [Wil21]. Analysis [NI21]. analytics [Che18]. Applications [HN17, Koz16, Str20, Wil21, Agg18, And19, Che18, Gue19, LK23, TG23, Tor17]. Approach [Bod21]. Apps [Var15, AH17]. Architecture [MKSS18]. arrgh [Kor18]. Art [Yoo21]. ARV [SC17]. asynchronous [SC17]. automate [DJ22, LK23]. Automation [LK23].

based [AH17, YKN^{+}22]. Basics [Yoo21]. Bazel [McN19]. beautiful [Wil21]. Beginners [Tho21, Val17]. Beginning [McN19]. bíogo [KPA17, KA17]. bíogo/hts [KPA17]. bíogo/ncbi [KA17]. Black [SPK20]. blueprints [Rye16]. Boost [Str20]. brain [McG19]. brain-friendly [McG19]. Breaking [CMAS22]. Build [Dox16, TG23, Agg18, And19, Gue19, LK23, Rye16, Tor17]. Building [HN17, Jac17, Koz16, McN19, Var15, Wil21, Tit21].

case [IHS16]. challenges [Gun18a]. clean [Sco18]. Cloud [AH17, HN17, Tit21, DJ22]. Code [Ric17, Woo21, Sco18]. Compiler [MKSS18, Bal19]. complete [Val17]. concurrent [Gue19]. Containerize [Ric17]. cookbook [Agg18, Tor17]. course [Val17]. Create [Wil21]. Cross [Wil21]. Cross-Platform [Wil21]. cutting [Rye16]. cutting-edge [Rye16]. CYPHONIC [YKN^{+}22].

Debt [NI21]. demonstrating [Che18]. dependency [Sco18]. Design [VCR17]. develop [AH17, Sco18]. Development [TG23, Var15, Agg18, And19]. DevOps [DJ22]. distributed [Tor19].

easier [Sco18]. edge [Rye16]. eight [Che18]. Empirical [CMAS22]. end [Che18]. end-to-end [Che18]. environment [CGP^{+}22]. environments [Tit21]. Essential [New17]. Evaluation [YKN^{+}22]. Experiential [NI21]. explore [Gun18a]. extended [KMMR23].

First [McG19]. framework [KMMR23]. friendly [McG19]. Full [TG23, Agg18, And19]. Full-Stack [TG23, Agg18]. functional [SC17]. Fyne [Wil21].

Game [CKN20]. Gap [NI21]. generation [IHS16]. Genius [Tho21]. Getting [Spa18]. gigabyte [Lem21]. Gin [And19]. GitHub [DJ22]. Go [And17a, BF16, Ket17, Whi17, Agg18, AH17, And19, Ano12, Bal17, Bal18, Bal19, Bod21, Cha16, Che18, CKN20, CMAS22, CGP^{+}22, Dan18, DJ22, DK16, Dox16, Fei22, Gue19, Gun18b, Gun18a, HN17, IHS16, Jac17, Ken16, Ken17, Ken18, KMMR23, KPA17, KA17, Kor18, Koz16, LK23, McG19, McN19, Mey14, MKSS18, NI21, New17, Ram18, Ric17, Rye16, SC17, Sco18, Spa18, SPK20, Str20, Tho21, Tit21, TG23, Tor17, Tor19, Val17, Var15, Viv16, VCR17, Wil21, Woo21, Yoo21, YKN^{+}22]. Go2Pins [KMMR23]. Golang [AH17, And17b, Gue19, Str20]. Google [Fei22, Val17]. GopherJS [And19]. graphical [Wil21]. GUI [Wil21]. Guide [Tho21, McG19, Viv16].

Hackers [SPK20]. handling [KPA17].

Hands [And19, Gue19, Gun18b, Gun18a, Sco18, Str20]. Hands-on [And19, Gue19, Gun18b, Gun18a, Sco18, Str20]. Head [McG19]. High [Str20, AH17, KPA17]. hts [KPA17].

Idiomatic [Bod21]. injection [Sco18]. insightful [Viv16]. interface [Kor18]. interfaces [KA17]. Interpreter [Bal18]. Introducing [Dox16, Yoo21]. Isomorphic [Bal17].

Java [Fei22, McN19].

Kubernetes [DJ22].

Language [Ano12, CKN20, DK16, Fei22, Mey14, MKSS18, Spa18, CGP^{+}22, DJ22, IHS16, KPA17, KA17, LK23, TG23, Viv16, Wil21, YKN^{+}22]. Languages [NI21]. Learn [DJ22, Fei22, LK23]. Learning [Bod21, Viv16, Whi17, Che18]. Linux [Gue19]. LTL [KMMR23].

Machine [Whi17, Che18]. maintain [Sco18]. Mastering [And17a]. method [IHS16]. microservice [AH17]. microservice-based [AH17]. Microservices [HN17, Jac17]. MIPS32 [MKSS18]. model [SC17]. Modeling [CKN20]. Modern [And17b, Gue19]. modular [Tor17].

Native [HN17, Tit21, AH17]. Natural [CKN20]. ncbi [KA17]. Network [LK23, New17, Woo21, YKN^{+}22]. Networks [New17]. Number [Lem21].

OpenCPU [Kor18]. operations [LK23]. Optimize [Str20]. organisation [SC17]. Overlay [YKN^{+}22].

Package [CMAS22]. Packer [DJ22]. parsing [Lem21]. Patterns [VCR17].

Pentesters [SPK20]. Performance [Str20, AH17]. Phrasebook [Ano12]. Platform [Wil21]. platform-agnostic [Wil21]. Play [CKN20]. practice [BF16]. predictive [Che18]. Procedural [NI21]. production [Rye16]. production-ready [Rye16]. Programmers [Fei22]. Programming [And17a, And17b, Ano12, Bod21, DK16, Fei22, Ken17, Ken18, Mey14, New17, Spa18, SPK20, Tho21, Woo21, Yoo21, AH17, Cha16, CGP^{+}22, Gue19, Gun18a, LK23, Rye16, TG23, Val17, Viv16, Wil21]. Programs [Dox16, KMMR23]. Projects [VCR17, Che18]. protocol [YKN^{+}22]. Python [Ram18].

quickly [TG23].

R [Kor18]. React [And19, HN17]. read [Sco18]. readable [Tor17]. ready [Rye16]. Real [Bod21, VCR17, Gun18a, Rye16]. Real-World [VCR17, Bod21, Gun18a, Rye16]. Reliable [Dox16, Woo21, Tit21]. Report [NI21]. Resilience [Str20]. RESTful [Var15]. RISC [SC17]. RISC-V [SC17].

Safety [CMAS22]. Scalable [Dox16, Var15]. Scale [Str20]. Scratch [Woo21]. second [Lem21]. Secure [Woo21]. Securing [New17]. Security [Dan18]. sequence [KPA17]. server [Kor18]. servers [DJ22]. Services [Var15, Woo21, KA17, Tit21]. Skills [New17]. solutions [Rye16, Tor19]. solving [Gun18a]. Stack [TG23, Agg18, And19]. started [Spa18]. Study [CMAS22]. Support [MKSS18]. system [Gue19, Kor18]. systems [Gue19, Tor19].

Technical [NI21]. techniques [Rye16]. technology [Rye16]. Terraform [DJ22]. Test [IHS16, Sco18]. testable [Tor17].

Testing [Ket17, McN19, Tor19]. throughput [KPA17]. Transformer [CKN20]. Type [CMAS22].

Ultimate [Ken17, Ken18]. Unix [Gue19]. unreliable [Tit21]. unsafe [CMAS22]. Usage [CMAS22]. use [DJ22]. Using [New17, Gue19, LK23, Rye16, TG23, Wil21].

V [SC17]. verification [KMMR23]. version [Bal18, KMMR23]. vs [Ram18]. Vue.js [TG23].

Web [TG23, Agg18, AH17, And19, Cha16, HN17, Koz16, Var15]. World [VCR17, Bod21, Gun18a, Rye16]. Writing [Bal18, Bal19].

References

Aggarwal:2018:GWD

[Agg18] Arpit Aggarwal. Go web development cookbook: build fullstack web applications with Go. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78728-674-6, 1-78728-856-0. viii + 321 pp. LCCN QA76.73.G63; QA76.76.A65 .A343 2018.

Andrawos:2017:CNP

[AH17] Mina Andrawos and Martin Helmich. Cloud native programming with Golang: develop microservice-based high performance web apps for the cloud with Go. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78712-598-X (print), 1-78712-796-6 (e-book). vi + 385 pp. LCCN QA76.76.D47; QA76.73.G63 .A537 2017. URL

http://proquestcombo.safaribooksonline.
com/9781787125988.

Andrawos:2017:MGP

[And17a] Mina Andrawos. Mastering Go Programming. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78646-823-9. 1 streaming video file, approximately 20 hr., 13 min. pp. LCCN ????

Andrawos:2017:MGPA

[And17b] Mina Andrawos. Modern Golang Programming. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78712-525-4. 1 streaming video file (14 hours, 37 minutes) pp. LCCN ????

Andrawos:2019:HFS

[And19] Mina Andrawos. Hands-on full stack development with Go: build full stack web applications with Go, React, Gin, and GopherJS. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78913-865-

5. v + 311 pp. LCCN QA76.73.G63. URL http://

proquest.safaribooksonline.
com/?fpi=9781789130751.

Anonymous:2012:GPL

[Ano12]  Anonymous.  The Go Pro-

gramming Language Phrasebook. Addison-Wesley Professional, ????, 2012. ISBN 0-13-291896-X. ???? pp.

Balasubramanian:2017:IG

[Bal17] Kamesh Balasubramanian. Isomorphic Go. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78839-418-6 (print). 505 (est.) pp. LCCN QA76.73.G63.

Ball:2018:WIG

[Bal18] Thorsten Ball. Writing an Interpreter in Go: version 1.6. Thorsten Ball, ????, 2018. ISBN 3-9820161-1-8 (paperback). 263 pp. LCCN QA76.76.C65 .B35 2018.

Ball:2019:WCG

[Bal19] Thorsten Ball. Writing a compiler in Go. Thorsten Ball, Aschaffenburg, Germany, version 1.1. edition, 2019. ISBN 3-9820161-0-X. 351 pp. LCCN QA76.73.G63 B348 2019.

Butcher:2016:GP

[BF16] Matt Butcher and Matt Farina. Go in practice. Manning Publications, Greenwich, CT, USA, 2016. ISBN 1-63343-007-3. xxi + 287 pp. LCCN QA76.73.G63 B87

                                        2016. URL http://proquest.
                                        tech.safaribooksonline.de/
                                        9781633430075.

Bodner:2021:LGI

[Bod21]  Jon Bodner.  Learning Go:

an Idiomatic Approach to Realworld Go Programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2021. ISBN 1-4920-7721-6. xvii + 352 pp. LCCN QA76.73.G63 B64 2021.

Cox:2022:GPL

[CGP^{+}22] Russ Cox, Robert Griesemer, Rob Pike, Ian Lance Taylor, and Ken Thompson. The Go programming language and environment. Communications of the ACM, 65(5):70–78, May 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/
10.1145/3488716.

Chang:2016:GWP

[Cha16] Sau Sheong Chang. Go Web programming. Manning Publications, Greenwich, CT, USA, 2016. ISBN 1-61729-256-7. xx + 292 pp. LCCN QA76.73.G63.

URL http://proquest.tech.
safaribooksonline.de/9781617292569.

Chew:2018:GML

[Che18]  Xuanyi Chew.  Go machine

learning projects: eight projects demonstrating end-to-end machine learning and predictive analytics applications in Go. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78899-340-3, 1-78899-519-8. 339 pp. LCCN QA76.73.G63.

Ciolino:2020:GTN

[CKN20] Matthew Ciolino, Josh Kalin, and David Noever. The Go transformer: Natural language modeling for game play. In IEEE, editor, 2020 Third International Conference on Artificial Intelligence for Industries (AI4I), pages 23–26. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020.

Costa:2022:BTS

[CMAS22] Diego Elias Costa, Suhaib Mujahid, Rabe Abdalkareem, and Emad Shihab. Breaking type safety in Go: An empirical study on the usage of the unsafe package. IEEE Transactions on Software Engineering, 48(7): 2277–2294, July 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Daniel:2018:SG

[Dan18] John Leon Daniel. Security with Go. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78862-791-1. LCCN QA76.585 .L466 2018; QA76.59.

Doak:2022:GDL

[DJ22] John Doak and David Justice. Go for DevOps: Learn how to use the Go language to automate servers, the cloud, Kubernetes, GitHub, Packer, and Terraform. Packt Publishing, Birmingham, UK, 2022. ISBN 1-80181-889-4 (paperback), 1-80181-934-3 (ebook). xxiv + 608 pp. LCCN QA76.73.G63 D63 2022.

Donovan:2016:GPL

[DK16] Alan A. A. Donovan and Brian W. Kernighan. The Go Programming Language. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 2016. ISBN 0-13-419044-0 (paperback). xvii + 380 pp. LCCN QA76.73.G63 D66 2016. URL https:/

/www.pearson.com/us/higher-
education/program/Donovan-
Go-Programming-Language-The/
PGM234922.html.

Doxsey:2016:IGB

[Dox16] Caleb Doxsey. Introducing Go:

Build Reliable, Scalable Programs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2016. ISBN 1-4783-5582-4, 1-4919-4195-2 (paperback). x + 111 pp. LCCN QA76.73.G63 D69

2016. URL http://proquest.
tech.safaribooksonline.de/
9781491941997.

Feigenbaum:2022:GJP

[Fei22]  Barry Feigenbaum.  Go for

Java Programmers Learn the Google Go Programming Language. Apress, Berkeley, CA, USA, 2022. ISBN 1-4842-7199-8. xxxv + 582 + 61 + 1 pp. LCCN QA76.73.G63 F45 2022.

Guerrieri:2019:HSP

[Gue19] Alex Guerrieri. Hands-on system programming with Go: build modern and concurrent applications for Unix and Linux systems using Golang. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78980-336-5 (e-book), 1-78980-407-8. xii + 437 pp. LCCN QA76.73.G63. URL http://

proquest.safaribooksonline.
com/?fpi=9781789804072.

Guney:2018:HGP

[Gun18a] Tarik Guney. Hands-on Go programming: explore Go by solving real-world challenges. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78953-487-9 (e-book). iii + 166 pp. LCCN QA76.73.G63 G86 2018.

Guney:2018:HG

[Gun18b] Tarik Guney. Hands-on with Go. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78913-251-7. 1 streaming video file (2 hours, 26 minutes) pp. LCCN ????

Hoffman:2017:CNG

[HN17] Kevin Hoffman and Dan Nemeth. Cloud Native Go: Building Web Applications and Microservices for the Cloud with Go and React. Addison-Wesley Professional, ????, 2017. ISBN 0-13-450575-1 (e-book), 0-13-450578-6, 0-13-450580-8, 0-672-33779-7. 256 pp. LCCN QA76.585. URL http:

                                        //www.vlebooks.com/vleweb/
                                        product/openreader?id=none\
                                        %26isbn=9780134505800.

Irawan:2016:TCG

[IHS16] Edbert Wijaya Irawan, Bayu Hendradjaya, and Wikan Danar Sunindyo. Test case generation method for Go language. In IEEE, editor, 2016 International Conference on Data and Software Engineering (ICoDSE), pages 1– 5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2016.

Jackson:2017:BMG

[Jac17] Nic Jackson. Building Microservices with Go. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78646-866-2, 1-78646-979-0. ix + 338 pp. LCCN QA76.73.G63.

Kortschak:2017:PBOb

[KA17] R. Daniel Kortschak and David L.

Adelson. bíogo/ncbi: interfaces to NCBI services for the Go language. Journal of Open Source Software, 2(18):234:1–234:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/10.
21105/joss.00234.

Kennedy:2016:GA

[Ken16] William Kennedy. Go in action. Manning Publications, Greenwich, CT, USA, 2016. ISBN 1-61729-178-1. xix + 241 pp. LCCN ????

Kennedy:2017:UGP

[Ken17] William Kennedy. Ultimate Go Programming. Addison-Wesley Professional, 2017. ISBN 0-13-475747-5. 1 video file (14 hours, 51 minutes) pp. LCCN ????

Kennedy:2018:UGP

[Ken18]  William Kennedy.  Ultimate

Go Programming. Addison-Wesley Professional, second edition, 2018. ISBN 0-13-526165-1. 1 video file (16 hours, 51 minutes) pp. LCCN ????

Ketelsen:2017:TG

[Ket17]  Brian Ketelsen.  Testing in

Go. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-7276-9, 1-4919-7277-7. 1 video file (25 minutes) pp. LCCN ????

Kirszenberg:2023:GFL

[KMMR23] Alexandre Kirszenberg, Antoine Martin, Hugo Moreau, and Etienne Renault. Go2Pins: a framework for the LTL verification of Go programs (extended version). International Journal on Software Tools for Technology Transfer (STTT), 25(1):77–94, February 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https:/

/link.springer.com/article/
10.1007/s10009-022-00692-w.

Kortschak:2018:PAG

[Kor18] R. Daniel Kortschak. arrgh: a Go interface to the OpenCPU R server system. Journal of Open Source Software, 3(21): 517:1, January 2018. CO-DEN ???? ISSN 2475-9066.

URL http://joss.theoj.org/
papers/10.21105/joss.00517.

Kozyra:2016:GBW

[Koz16] Nathan Kozyra. Go: Building Web Applications (1). Packt Publishing, Birmingham, UK, 2016. ISBN 1-78712-349-9, 1-78712-659-5 (e-book). 665 pp. LCCN TK5105.8885.L674.

URL http://www.myilibrary.
com?id=952170.

Kortschak:2017:PBOa

[KPA17] R. Daniel Kortschak, Brent S. Pedersen, and David L. Adelson. bíogo/hts: high throughput sequence handling for the Go language. Journal of Open Source Software, 2(10):168:1, February 2017. CODEN ????

ISSN 2475-9066. URL http:

//joss.theoj.org/papers/10.
21105/joss.00168.

Lemire:2021:NPG

[Lem21] Daniel Lemire. Number parsing at a gigabyte per second. Software —Practice and Experience, 51(8): 1700–1727, August 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leiva:2023:NAG

[LK23] Nicolas Leiva and Michael Kashin.

Network Automation with Go: Learn how to automate network operations and build applications using the Go programming language. Packt Publishing, Birmingham, UK, 2023. ISBN 1-80056-092-3 (paperback), 1-80056-101-6 (e-book). xviii + 423 pp. LCCN QA76.73.G63.

McGavren:2019:HFG

[McG19] Jay McGavren. Head First Go:

a brain-friendly guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4919-6955-5. xxxiv + 520 pp. LCCN QA76.73.G63 M34 2019.

McNerney:2019:BBB

[McN19] P. J. McNerney. Beginning Bazel:

Building and Testing for Java, Go, and More. Apress, Berkeley, CA, USA, 2019. ISBN 1-4842-5194-6. xi + 192 pp. LCCN QA75.5-76.95. URL http:

//www.vlebooks.com/vleweb/
product/openreader?id=none\
%26isbn=9781484251942.

Meyerson:2014:GPL

[Mey14] Jeff Meyerson. The Go programming language. IEEE Software, 31(5):104, September/October 2014. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). URL http://csdl.

                                        computer.org/csdl/mags/so/
                                        2014/05/mso2014050104.html.

Mladenovi:2018:ASM

[MKSS18] Dragan Mladenoviç, Milan Knežević, Vladimir Stefanović, and Miloš Subotić. Adding support for the MIPS32 architecture into the Go language compiler. In IEEE, editor, 2018 26th Telecommunications Forum (TELFOR), pages 420–425. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2018.

Newmarch:2017:NPG

[New17] Jan Newmarch. Network Programming with Go: Essential Skills for Using and Securing Networks. Apress, Berkeley, CA, USA, 2017. ISBN 1-4842-2691-7, 1-4842-2692-5. xxi + 274 + 36 pp. LCCN QA76.73.G63 N48 2017; QA76.76.C65.

Nelson:2021:GAT

[NI21] Grant Nelson and Clemente Izurieta. A gap in the analysis of technical debt in procedural languages: an experiential report on Go. IEEE Software, 38(6): 71–75, 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ramalho:2018:PVG

[Ram18] Luciano Ramalho. Python vs.

Go. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2018. ISBN 1-4920-3317-0, 1-4920-3318-9. ???? pp. LCCN QA76.73.P98.

Rice:2017:HCY

[Ric17] Liz Rice. How to Containerize Your Go Code. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-8230-6. ???? pp. LCCN ????

Ryer:2016:GPB

[Rye16] Mat Ryer. Go programming blueprints: build real-world, production-ready solutions in Go using cutting-edge technology and techniques. Packt Publishing, Birmingham, UK, second edition, 2016. ISBN 1-78646-247-8, 1-78646-894-8. viii + 376 pp. LCCN ????

Sartori:2017:GFM

[SC17] Marcos L. L. Sartori and Ney L. V. Calazans. Go functional model for a RISC-V asynchronous organisation — ARV. In IEEE, editor, 2017 24th IEEE International Conference on Electronics, Circuits and Systems (ICECS), pages 381–348. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017.

Scott:2018:HDI

[Sco18] Corey Scott. Hands-on dependency injection in Go: develop clean Go code that is easier to read, maintain, and test. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78913-175-8. 338 pp. LCCN QA76.76.A65 .S368 2018; QA76.73.G63. URL http:/

/proquest.safaribooksonline.
com/?fpi=9781789132762.

Spaulding:2018:GSG

[Spa18] Matthew Spaulding. Getting started with Go Programming Language. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78847-185-7. 1 streaming video file (3 hours, 4 minutes) pp. LCCN ????

Steele:2020:BGG

[SPK20] Tom Steele, Chris Patten, and Dan Kottmann. Black Hat Go: Go Programming for Hackers and Pentesters. No Starch Press, San Francisco, CA, USA, 2020. ISBN 1-0981-2264-X, 1-59327-865-9. xxv + 337 pp. LCCN QA76.9.A25 S739 2020.

Strecansky:2020:HHP

[Str20]  Bob Strecansky.  Hands-on

High Performance with Go: Boost and Optimize the Performance of Your Golang Applications at Scale with Resilience. Packt Publishing, Birmingham, UK, 2020. ISBN 1-78980-177-X (paperback), 1-78980-578-3. vii + 391 pp. LCCN QA76.73.G63. URL http://

proquest.safaribooksonline.
com/?fpi=9781789805789.

Tolaram:2023:FSW

[TG23] Nanik Tolaram and Nick Glynn.

Full-Stack Web Development with Go: Build your web applications quickly using the Go programming language and Vue.js. Packt Publishing, Birmingham, UK, 2023. ISBN 1-80323-419-9 (paperback), 1-80324-391-0 (e-book). xvii + 282 pp. LCCN QA76.73.G63.

Thornton:2021:GBG

[Tho21] Edward Thornton. Go for Beginners: A Genius Guide to Go Programming. ????, ????, 2021. ISBN ???? ???? pp. LCCN ????

Titmus:2021:CNG

[Tit21] Matthew Titmus. Cloud Native Go: building reliable services in unreliable environments. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2021. ISBN 1-4920-7633-3. ???? pp. LCCN ????

Torres:2017:GCB

[Tor17] Aaron Torres. Go cookbook: build modular, readable, and testable applications in Go. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78328-684-9. 392 (est.) pp. LCCN QA76.73.G63.

Torres:2019:ASG

[Tor19] Aaron Torres. Advanced solutions in Go testing and distributed systems. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78862-788-1. 1 streaming video file (1 hour, 48 minutes) pp. LCCN ????

Valentine:2017:CGG

[Val17] DaveIn Valentine. The complete Google Go programming course for beginners. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78862-697-4. 1 streaming video file (4 hours, 50 minutes) pp. LCCN ????

Varghese:2015:WDG

[Var15] Shiju Varghese. Web Development with Go: Building Scalable Web Apps and RESTful Services. Apress, Berkeley, CA, USA, 2015. ISBN 1-4842-1052-2. xx + 289 + 69 pp. LCCN ????

Vivien:2017:GDP

[VCR17] Vladimir Vivien, Mario Contreras, and Mat Ryer. Go: Design Patterns for Real-World Projects. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78839-055-5, 1-78839-287-6. 1091 pp. LCCN ????

Vivien:2016:LGPh

[Viv16] Vladimir Vivien. Learning Go programming: an insightful guide to learning the Go programming language. Packt Publishing, Birmingham, UK, 2016. ISBN 1-78439-233-2, 1-78439-543-9. ix + 327 pp. LCCN QA76.73.G63.

URL http://proquest.tech.
safaribooksonline.de/9781784395438.

Whitenack:2017:MLG

[Whi17] Daniel Whitenack. Machine Learning With Go. Packt Publishing; O’Reilly Media Inc., [Erscheinungsort nicht ermittelbar]; Sebastopol, CA, 2017. ISBN 1-78588-210-4, 1-78588-390-9. 287 (??) pp. LCCN QA76.73.G63 W55 2017.

Williams:2021:BCP

[Wil21]  Andrew Williams.  Building

Cross-Platform GUI Applications with Fyne: Create beautiful, platform-agnostic graphical applications using Fyne and the Go programming language. Packt Publishing, Birmingham, UK, 2021. ISBN 1-80056-316-7 (paperback), 1-80056-688-3 (PDF ebook). xii + 301 pp. LCCN QA76.9.U83 .W555 2021.

Woodbeck:2021:NPG

[Woo21] Adam Woodbeck. Network Programming with Go: Code Secure and Reliable Network Services from Scratch. No Starch Press, San Francisco, CA, USA, 2021. ISBN 1-0981-2889-3, 1-71850-088-2 (paperback). xxii + 363 pp. LCCN QA76.73.G63 W663 2021.

Yoshikawa:2022:ENC

[YKN^{+}22] Taiki Yoshikawa, Hijiri Komura, Chihiro Nishiwaki, Ren Goto, Kazushige Matama, and Katsuhiro Naito. Evaluation of new CYPHONIC: Overlay network protocol based on Go language. In IEEE, editor, 2022 IEEE International Conference on Consumer Electronics (ICCE), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Yoon:2021:AGB

[Yoo21] Harry Yoon. The Art of Go — Basics: Introducing to Programming in Go. Coding Books Press, ????, 2021. ???? pp. LCCN ????